### v1

# Graph Alignment via Dual-Pass Spectral Encoding and Latent Space Communication

Graph Alignment via Dual-Pass Spectral Encoding and Latent Space Communication
Maysam Behmanesh
LIX, École Polytechnique, IP Paris
&Erkan Turan
LIX, École Polytechnique, IP Paris
Maks Ovsjanikov
LIX, École Polytechnique, IP Paris
Graph alignment—the problem of identifying corresponding nodes across multiple graphs—is fundamental to numerous applications. Most existing unsupervised methods embed node features into latent representations to enable cross-graph comparison without ground-truth correspondences. However, these methods suffer from two critical limitations: the degradation of node distinctiveness due to oversmoothing in GNN-based embeddings, and the misalignment of latent spaces across graphs caused by structural noise, feature heterogeneity, and training instability, ultimately leading to unreliable node correspondences. We propose a novel graph alignment framework that simultaneously enhances node distinctiveness and enforces geometric consistency across latent spaces. Our approach introduces a dual-pass encoder that combines low-pass and high-pass spectral filters to generate embeddings that are both structure-aware and highly discriminative. To address latent space misalignment, we incorporate a geometry-aware functional map module that learns bijective and isometric transformations between graph embeddings, ensuring consistent geometric relationships across different representations. Extensive experiments on graph benchmarks demonstrate that our method consistently outperforms existing unsupervised alignment baselines, exhibiting superior robustness to structural inconsistencies and challenging alignment scenarios. Additionally, comprehensive evaluation on vision-language benchmarks using diverse pretrained models shows that our framework effectively generalizes beyond graph domains, enabling unsupervised alignment of vision and language representations.
1 Introduction
Graph alignment—also referred to as network alignment or graph matching—is a fundamental problem in machine learning and graph theory, concerned with identifying a correspondence between the nodes of two graphs such that structurally similar or semantically equivalent nodes are matched.
Graph alignment arises in a wide range of application domains, including bioinformatics (e.g., protein interaction networks) (Liao et al., 2009; Singh et al., 2007), social network analysis (Li et al., 2018; Korula and Lattanzi, 2014), computer vision (Liu et al., 2022a; Chen et al., 2025; Wang et al., 2019), and natural language processing (Osman and Barukub, 2020; Guillaume, 2021). Due to its combinatorial nature, graph alignment is computationally challenging, often requiring approximation or heuristic algorithms.
Graph alignment methods are typically classified into three categories based on their alignment strategies: optimization-based, optimal transport–based, and embedding-based approaches. They also vary in the level of supervision required, ranging from unsupervised to semi-supervised—using partial node correspondences—and fully supervised methods. A detailed overview of these categories with related works is provided in Appendix A.
Embedding-based graph alignment methods encode graphs into low-dimensional node representations via Graph Neural Networks (GNNs) (He et al., 2024; Fey et al., 2020; Gao et al., 2021a), followed by alignment through transformations or joint learning with cross-graph regularization. Node matching is then performed using nearest-neighbor search or assignment algorithms, achieving better scalability than optimization-based alternatives.
These methods are typically formulated as unsupervised learning tasks, where ground-truth node correspondences across graphs are unavailable. Despite their computational advantages, these approaches face several inherent challenges that limit their effectiveness and reliability:
1) Degradation of node distinctiveness in GNN embeddings. While GNNs generate node representations by aggregating features from local neighborhoods to capture structural information, this aggregation process inherently reduces node distinctiveness. This limitation is particularly problematic for graph alignment, where accurate correspondence identification depends on highly discriminative representations. As embeddings lose uniqueness, alignment becomes increasingly ambiguous and error-prone.
2) Misaligned latent spaces across graphs. In the absence of supervision, explicit constraints, or alignment-specific objectives during training, embedding-based methods struggle to produce comparable latent spaces across different graphs. Even when using shared encoders, the resulting embeddings often occupy misaligned geometric spaces due to structural inconsistencies, feature heterogeneity, and training instability. As illustrated in Figure 1, nodes with identical local structures may be mapped to distant regions in their respective latent spaces.
This misalignment arises from multiple sources. First, structural inconsistencies—such as missing or noisy edges—distort neighborhood aggregation during message passing, leading to incompatible embeddings for otherwise corresponding nodes. Second, feature inconsistency across graphs—stemming from differences in user attributes, schema, or data domains—causes graph encoders to embed semantically equivalent nodes into disjoint subspaces. Lastly, GNN-based encoders often exhibit stochasticity in training; different random initializations can yield drastically different embeddings, even on fixed graph inputs (Moschella et al., 2023). Without explicit mechanisms to harmonize or align the latent spaces, these inconsistencies severely hinder cross-graph communication and undermine the reliability of node alignment.
Figure 1 illustrates the latent spaces learned by a 2-layer GCN on synthetic graphs. Panel (a) shows that identical graphs produce well-aligned embeddings, facilitating effective correspondence detection. However, panels (b) and (c) reveal that minor feature and structural inconsistencies cause corresponding node embeddings to diverge significantly, compromising alignment quality. Most critically, panel (d) shows that retraining the same model on identical graphs with different random initializations produces drastically different latent spaces, underscoring the instability and non-deterministic nature of learned representations.
Figure 1: Illustration of limitations in embedding-based graph alignment using synthetic data. $\mathcal{G}_{1}$ is a ring graph with 100 nodes and 2D random features. (a) $\mathcal{G}_{2}$ is identical to $\mathcal{G}_{1}$, yielding well-aligned embeddings. (b) Feature inconsistency introduced through Gaussian noise (std=0.2) causes divergence. (c) Structural inconsistency via 30% edge dropout distorts the embedding alignment. (d) Both graphs are identical, but different training runs produce inconsistent embedding spaces, highlighting instability in unsupervised latent space learning.
In this paper, we introduce GADL—Graph Alignment with Dual-pass encoder and Latent space communication—which builds upon the Graph Autoencoder (GAE) framework (Kipf and Welling, 2016) by incorporating a dual-pass encoding architecture and cross-graph latent communication mechanism tailored for unsupervised graph alignment tasks.
First, to address the degradation of node distinctiveness caused by oversmoothing in GNN neighborhood aggregation, GADL employs a dual-pass GCN encoder that combines low-pass and high-pass spectral filters. The low-pass branch captures structural context, while the high-pass branch preserves fine-grained node distinctiveness. Their concatenation yields embeddings that are both structure-aware and highly discriminative—crucial for accurate graph alignment. Second, to address latent space misalignment across graphs, GADL incorporates a geometry-aware functional map module that learns explicit transformations between different graph embeddings. By enforcing bijectivity and orthogonality constraints, this module ensures that embeddings across graphs become mutually consistent and locally isometric, enabling effective cross-graph communication and alignment without requiring ground-truth node pairs.
Our key contributions are summarized as follows:
  1. 1.
    We propose a novel dual-pass GCN encoder that combines low-pass and high-pass spectral filters to produce embeddings that are both structure-aware and highly discriminative.
  2. 2.
    We introduce a geometry-aware functional map module that explicitly aligns latent spaces across graphs, enabling robust cross-graph communication without supervision.
  3. 3.
    We conduct extensive experiments on graph alignment benchmarks, demonstrating superior performance and robustness to structural inconsistencies in unsupervised alignment tasks.
  4. 4.
    We evaluate our framework on vision-language benchmarks, demonstrating that our framework effectively generalizes beyond graph domains to enable cross-modal alignment.
2 Preliminaries
We formaly define the problem of aligning attributed nodes from a source graph $\mathcal{G}_{s}$ to a target graph $\mathcal{G}_{t}$ in an unsupervised setting. The goal is to identify, for each node in the source graph, a corresponding node in the target graph.
Definition 1 (Graph Alignment (GA)).
Given two graphs $\mathcal{G}_{s}=(\mathcal{V}_{s},\mathcal{E}_{s},\mathbf{X}_{s})$ and $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t})$, where $\mathcal{V}$ denotes the set of nodes, $\mathcal{E}$ the set of edges, and $\mathbf{X_{*}}\in\mathbb{R}^{N_{*}\times k_{*}}$ the associated node attributes (features), the graph alignment problem aims to find a one-to-one mapping $\pi:\mathcal{V}_{s}\rightarrow\mathcal{V}_{t}$ such that for each node $u\in\mathcal{V}_{s}$, $\pi(u)=v\in\mathcal{V}_{t}$ and $\pi^{-1}(v)=u$. The objective is to identify correspondences between nodes in $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ that preserve structural similarity and attribute consistency across the two graphs.
We assume the GA problem between two general graphs with different number of nodes ($|\mathcal{V}_{s}|\neq|\mathcal{V}_{t}|$) in an unsupervised setting, where no ground-truth node correspondences are available during training, and the alignment depends solely on the structural and attribute information of the graphs.
2.1 Graph autoencoder for unsupervised node embedding
Graph autoencoders (GAEs) (Kipf and Welling, 2016) are graph representation learning frameworks used to learn node embeddings in an unsupervised setting. They aim to generate low-dimensional representations that effectively capture both node features and the underlying geometric structure of the graph.
Following the general principle of autoencoders, a GAE consists of two main components: an encoder $q_{\theta}(\mathbf{Z}\mid\mathcal{G})$ that maps the input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$—where $\mathbf{X}\in\mathbb{R}^{|V|\times k}$—into a latent embedding matrix $\mathbf{Z}\in\mathbb{R}^{|V|\times d}$, leveraging both graph structure and node features to learn meaningful representations; and a decoder $p_{\phi}(\mathcal{G}\mid\mathbf{Z})$ that reconstructs the original graph structure and node attributes from these latent embeddings, producing an approximation $\hat{\mathcal{G}}$ of the input graph. The model is trained to minimize a loss function composed of a reconstruction loss $\mathcal{L}_{\text{rec}}$, which measures the difference between $\mathcal{G}$ and $\hat{\mathcal{G}}$, and optionally a regularization term $\mathcal{L}_{\text{reg}}$ on the latent space:
$\mathcal{L}=\mathcal{L}_{\text{rec}}(\mathcal{G},\hat{\mathcal{G}})+\lambda\mathcal{L}_{\text{reg}}(\mathbf{Z}),$
where $\lambda$ controls the strength of regularization. This framework enables unsupervised learning of node embeddings that capture the intrinsic geometric structure of the graph, thereby facilitating downstream tasks such as graph alignment.
In this framework, the encoder is typically implemented using a GNN $\phi(\mathbf{X},\mathbf{S};\theta):\mathbb{R}^{N\times k}\rightarrow\mathbb{R}^{N\times d}$ with parameters $\theta$, which maps node features $\mathbf{X}$ and graph structure $\mathbf{S}$ (e.g., an adjacency or normalized Laplacian matrix) to latent node embeddings $\mathbf{Z}$. The decoder is typically a simple, non-parametric function that reconstructs the graph structure from the learned embeddings. A common choice is the inner product decoder, which estimates the adjacency matrix $\hat{\mathbf{A}}$ as $\hat{\mathbf{A}}=\mathbf{Z}\mathbf{Z}^{\top}$. This formulation assumes that the similarity between node embeddings reflects the likelihood of an edge, enabling the reconstruction of the graph topology directly from the embedding space.
2.2 Functional map on graphs
The functional map framework, originally proposed for 3D shape correspondence (Ovsjanikov et al., 2012), offers a compact and flexible approach that converts the problem of finding a complex node-to-node correspondence into learning a small, low-dimensional operator $C$ that aligns functions represented in a spectral basis. This paradigm naturally extends to graphs (Fumero et al., 2025; Behmanesh et al., 2024), where functions are defined on nodes, providing a powerful framework for comparing and aligning graph-structured data.
Building on the general framework of Deep Geometric Functional Maps (Donati et al., 2020), the functional map formulation is adapted to operate on graph-based latent representations through the following steps:
1. Feature extraction.
Given a pair of graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, each domain is associated with a set of descriptor functions, denoted by $\mathcal{F}_{\theta}({\mathcal{G}1})$ and $\mathcal{F}_{\theta}({\mathcal{G}_{2}})$, respectively. A descriptor function is a real-valued function defined on the nodes of a graph—either hand-crafted to capture structural information shared across graphs or learned via neural encoders—producing row feature matrices $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$.
2. Projection to spectral domain:
For each domain, the spectral basis $\Phi_{*}$ is computed by performing eigendecomposition of the normalized graph Laplacian, defined as $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. The descriptor functions are then projected onto the reduced spectral subspace $\Phi_{*}\in\mathbb{R}^{n_{*}\times r}$, spanned by the first $r$ eigenvectors, resulting in the spectral coefficients $\hat{\mathbf{F}}_{1}=\Phi_{1}^{\top}\mathbf{F}_{1}$, and $\hat{\mathbf{F}}_{2}=\Phi_{2}^{\top}\mathbf{F}_{2}$.
3. Functional map estimation.
A functional map $\mathbf{C}_{12}\in\mathbb{R}^{r\times r}$ is then estimated by aligning the spectral descriptors between the two domains via the following regularized least squares objective:
$\mathbf{C}_{12}=\arg\min_{\mathbf{C}}\|\mathbf{C}\hat{\mathbf{F}}_{1}-\hat{\mathbf{F}}_{2}\|_{F}^{2}+\alpha\|\Lambda_{2}\mathbf{C}-\mathbf{C}\Lambda_{1}\|_{F}^{2},$ (1)
where the second term is the Laplacian commutativity regularizer, which enforces that the functional map $\mathbf{C}_{12}$ approximately commutes with the graph Laplacians, thereby ensuring that $\mathbf{C}_{12}$ preserves their spectral properties.
3 Method overview
In this section, we introduce the proposed learning-based framework for the graph alignment problem. First, we present an overview of the framework, followed by a discussion of the inherent limitations of learning-based methods in graph alignment problems. Then, we present the motivation and the key innovations in the proposed framework that address these challenges. Figure 2 provides a schematic overview of the proposed framework.
Figure 2: Overview of the proposed framework. Given a pair of input graphs, the model uses a dual-pass encoder with shared parameters to extract node embeddings. To resolve latent space misalignment, a regularized functional map module is introduced to explicitly enforce structural constraints and enable communication between latent spaces. A graph decoder reconstructs the inputs, and the model is optimized with an overall loss. Finally, alignments are estimated via cosine similarity between the embeddings, followed by greedy matching to establish one-to-one correspondences.
3.1 Overall framework
Given two graphs $\mathcal{G}_{s}=(\mathcal{V}_{s},\mathcal{E}_{s},\mathbf{X}_{s})$ and $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t})$ as input, the objective is to learn meaningful latent representations of the nodes. The proposed architecture consists of two graph autoencoders with shared parameters, each designed to extract latent embeddings $\mathbf{Z}_{s}\in\mathbb{R}^{|V_{s}|\times d}$ and $\mathbf{Z}_{t}\in\mathbb{R}^{|V_{t}|\times d}$ from the respective input graphs. This is achieved by jointly encoding both the graph structure and node attributes through graph encoders $\mathbf{Z}_{s}=f^{s}_{\theta}(\mathbf{X}_{s},\mathbf{A}_{s})\in\mathbb{R}^{|V_{s}|\times d}$ and $\mathbf{Z_{t}}=f^{t}_{\theta}(\mathbf{X}_{t},\mathbf{A}_{t})\in\mathbb{R}^{|V_{t}|\times d}$, where $\mathbf{A}$ denotes the adjacency matrix encoding the graph structure, and $\mathbf{X}$ contains the node features.
As outlined in the introduction, this framework encounters two fundamental limitations that substantially reduce its effectiveness for graph alignment: loss of node distinctiveness and misaligned latent spaces across graphs in unsupervised settings. In the following sections, we present our proposed framework, which is specifically designed to address these limitations.
3.2 Graph encoder
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, a graph encoder $\mathbf{Z}=f_{\theta}(\mathbf{X},\mathbf{A})\in\mathbb{R}^{|V|\times d}$ embeds each node $v_{i}\in\mathcal{V}$ into a low-dimensional latent vector $\mathbf{z}_{i}\in\mathbb{R}^{d}$, such that the embeddings of neighboring nodes are encouraged to be similar. While this property allows the encoder to capture the local graph structure effectively, it poses a significant limitation for graph alignment tasks by reducing the distinctiveness of individual nodes—an essential factor for accurately identifying corresponding nodes across graphs.
Definition 2 (Ideal node embedding for graph alignment).
An ideal node embedding for graph alignment simultaneously achieves two key properties: local consistency, where the embeddings of neighboring nodes are similar such that $\max_{v\in V}\max_{u\in\mathcal{N}(v)}\|h_{v}^{(k)}-h_{u}^{(k)}\|$ is small, and global distinctiveness, where embeddings of distinct nodes are sufficiently different to allow unique identification, i.e., $\min_{v,w\in V}\|h_{v}^{(k)}-h_{w}^{(k)}\|$. As a result, accurate graph alignment requires embedding nodes that balance a fundamental trade-off: they must preserve local similarities to capture the underlying graph structure, while simultaneously preserving node distinctiveness by limiting excessive feature mixing to ensure that each node remains uniquely identifiable.
One of the simple yet effective graph encoders is the Graph Convolutional Network (GCN) (Kipf and Welling, 2017), which extends the convolution operation to graph-structured data. GCNs learn node representations by aggregating information from neighboring nodes, thereby capturing both node fetures and graph structure. A single GCN layer is defined by $\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right)$, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with added self-loops, $\tilde{\mathbf{D}}$ is the diagonal degree matrix of $\tilde{\mathbf{A}}$, $\mathbf{H}^{(l)}$ is the node feature matrix at the $l$-th layer, $\mathbf{W}^{(l)}$ is the learnable weight matrix of layer $l$, and $\sigma(\cdot)$ is a non-linear activation function such as ReLU.
Spectral interpretation of GCN: In graph signal processing, the graph Laplacian is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the degree matrix and $\mathbf{A}$ is the adjacency matrix. The Laplacian can be decomposed as $\mathbf{L}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}=(\mathbf{u}_{1},\ldots,\mathbf{u}_{n})$ is the matrix of eigenvectors, and $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{n})$ is the diagonal matrix of eigenvalues. The normalized graph Laplacian is defined as $\mathbf{L}_{\text{sym}}=\mathbf{D}^{-1/2}\mathbf{L}\mathbf{D}^{-1/2}$, whose eigenvalues $\lambda_{i}$ lie within the interval $[0,2]$.
The GCN filter can be expressed as $\tilde{\mathbf{A}}_{\text{GCN},\text{sym}}=\mathbf{I}-\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}(\mathbf{I}-\tilde{\mathbf{\Lambda}})\mathbf{U}^{\top}$, with an associated frequency response function $p_{\text{GCN}}(\tilde{\lambda}_{i})=1-\tilde{\lambda}_{i}$. Since the eigenvalues satisfy $\tilde{\lambda}_{i}\in[0,2)$, the response function $p_{\text{GCN}}(\tilde{\lambda}_{i})$ decreases as $\tilde{\lambda}_{i}$ increases, particularly over the range $[0,1]$. This behavior implies that the GCN filter primarily suppresses high-frequency components and thus acts as a low-pass filter in that region. However, for $\tilde{\lambda}_{i}>1$, $p_{\text{GCN}}(\tilde{\lambda}_{i})$ becomes negative, introducing noise and disrupting smoothness. This means GCN is not a completely low-pass filter and can degrade performance due to this issue.
Node embedding via spectral filtering: Low-pass filters preserve low-frequency components (small $\lambda_{i}$) and suppress high-frequency components, producing smooth embeddings where neighboring nodes have similar representations. Such embeddings are effective at capturing local structure and community information within the graph. In contrast, high-pass filters preserve high-frequency components (large $\lambda_{i}$), emphasizing the differences between neighboring nodes. This leads to embeddings that capture distinctive, discriminative features, making the latent representations of nodes more distinct and farther apart from those of their neighbors.
Dual-pass GCN encoder with spectral filtering: In our proposed model, we design a dual-encoder architecture comprising two complementary GCN variants that exploit the spectral properties of graph signals. The architecture consists of: 1) a low-pass GCN encoder $\mathbf{Z}_{l}=f_{\theta_{l}}(\mathbf{X},\mathbf{A})$ that aggregates information from neighboring nodes, and 2) a high-pass GCN encoder $\mathbf{Z}_{h}=f_{\theta_{h}}(\mathbf{X},\mathbf{A})$, which highlights differences between a node and its neighbors, generating distinctive embeddings that effectively capture discriminative features. The final node representation is obtained through concatenation $\mathbf{Z}=[\mathbf{Z}_{l}\,\|\,\mathbf{Z}_{h}]\in\mathbb{R}^{|V|\times(d_{l}+d_{h})}$.
Both encoders employ a unified spectral convolution framework with layer-wise propagation rule:
$\mathbf{Z}^{(m+1)}_{*}=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}_{*}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{Z}^{(m)}_{*}\mathbf{W}^{(m)}_{*}\right)$ (2)
[…]
The low-pass graph filter is characterized by $\tilde{\mathbf{A}}_{l,\text{sym}}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}_{l}\tilde{\mathbf{D}}^{-1/2}=\mathbf{I}-\frac{1}{2}\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}\left(\mathbf{I}-\frac{1}{2}\tilde{\mathbf{\Lambda}}\right)\mathbf{U}^{\top}$. This formulation reveals that $\tilde{\mathbf{A}}_{l,\text{sym}}$ exhibits a frequency response $p_{l}(\tilde{\lambda}_{i})=1-\frac{1}{2}\tilde{\lambda}_{i}$. The response function is monotonically decreasing over $\tilde{\lambda}_{i}\in[0,2]$, thereby attenuating high-frequency components while preserving smooth graph signals. This design enables the encoder to capture local structural patterns and maintain graph regularity in the learned embeddings. Similarly, the high-pass graph filter is defined by $\tilde{\mathbf{A}}_{h,\text{sym}}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}_{h}\tilde{\mathbf{D}}^{-1/2}=\frac{1}{2}\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}\left(\frac{1}{2}\tilde{\mathbf{\Lambda}}\right)\mathbf{U}^{\top}.$ This formulation indicates that $\tilde{\mathbf{A}}_{h,\text{sym}}$ acts as a spectral filter with frequency response $p_{h}(\tilde{\lambda}_{i})=\frac{1}{2}\tilde{\lambda}_{i}$. The frequency response $p_{h}(\tilde{\lambda}_{i})$ is a non-negative, monotonically increasing function over the eigenvalue range $\tilde{\lambda}_{i}\in[0,2]$. Consequently, the filter suppresses low-frequency components while amplifying high-frequency components, thereby functioning as a purely high-pass filter (see (Wang et al., 2022a) for details on the frequency response).
Theorem 1 (Discriminativity of dual-pass GCN encoder).
Let $\mathbf{z}_{i}^{\text{low}}\in\mathbb{R}^{d_{1}}$ and $\mathbf{z}_{i}^{\text{high}}\in\mathbb{R}^{d_{2}}$ denote the node embeddings of node $i$ in graph $\mathcal{G}$ produced by low-pass and high-pass GCN encoders, respectively and dual-pass embedding is defined as the concatenation $\mathbf{z}_{i}=[\mathbf{z}_{i}^{\text{low}}\,\|\,\mathbf{z}_{i}^{\text{high}}]\in\mathbb{R}^{d_{1}+d_{2}}$. Assume both input graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ use this encoder architecture to compute embeddings $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$, respectively. Then, the dual-pass GCN encoder constitutes an ideal node embedding for graph alignment by satisfying the following properties:
  1. 1.
    Spectral locality preservation: the embedding $\mathbf{z}_{i}$ preserves neighborhood similarity comparably to $\mathbf{z}_{i}^{\text{low}}$.
  2. 2.
    Enhanced node discriminability: the embedding $\mathbf{z}_{i}$ provides superior node correspondence discrimination compared to either component alone.
The proof is provided in Appendix B.
3.3 Latent space communication
While each GAE independently produces a latent space for its respective graph—resulting in misaligned embeddings—we address this limitation by incorporating deep functional maps to learn explicit mappings between latent representations. Rather than directly comparing raw embeddings—which may differ by arbitrary isometric transformations—we learn functional maps $\mathbf{C}^{12}$ and $\mathbf{C}^{21}$ that transform functions between latent spaces. These maps are optimized within our network using Equation (1), where $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ represent embeddings from the shared dual-pass encoder.
To facilitate latent space communication, our framework leverages spectral geometry principles and a regularized functional map module that enforces structural constraints. Specifically, we impose bijectivity and orthogonality losses to ensure the maps $\mathbf{C}^{12}$ and $\mathbf{C}^{21}$ are approximately invertible and locally isometric, preserving essential geometric properties.
The bijectivity loss promotes invertibility by ensuring that functions mapped from one latent space to the other and back are reconstructed accurately, thereby enforcing structural consistency and encouraging the learned feature spaces to be approximately invertible and mutually aligned through the learned maps. Formally, it is defined as:
$\mathcal{L}_{\text{bij}}=\|\mathbf{C}_{12}\mathbf{C}_{21}-\mathbf{I}\|_{F}^{2}+\|\mathbf{C}_{21}\mathbf{C}_{12}-\mathbf{I}\|_{F}^{2}.$ (3)
The orthogonality loss enforces that functional maps behave as partial isometries, preserving local geometry and structural information during cross-space transformations. This loss is given by:
$\mathcal{L}_{\text{orth}}=\|\mathbf{C}_{12}\mathbf{C}_{12}^{\top}-\mathbf{I}\|_{F}^{2}+\|\mathbf{C}_{21}^{\top}\mathbf{C}_{21}-\mathbf{I}\|_{F}^{2}.$ (4)
These regularizations enable geometry-aware alignment of latent spaces, facilitating reliable cross-graph alignment without requiring any ground-truth correspondences and effectively bridging independently learned embeddings.
3.4 Graph decoder
Given latent node embeddings $\mathbf{Z}=f_{\theta}(\mathbf{X},\mathbf{A})\in\mathbb{R}^{|V|\times d}$ produced by the encoder, the decoder reconstructs graph structure using inner product operations $\hat{\mathbf{A}}_{s}=\sigma(\mathbf{Z}_{s}\mathbf{Z}_{s}^{\top})$, and $\hat{\mathbf{A}}_{t}=\sigma(\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top})$, where $\sigma(\cdot)$ denotes the element-wise sigmoid function.
3.5 Model optimization and training Loss
Our model jointly optimizes GAE parameters and functional maps $\mathbf{C}_{12}$, $\mathbf{C}_{21}$ through end-to-end training. Given embeddings $\mathbf{Z}_{1}=f_{\theta}(\mathbf{X}_{s},\mathbf{A}_{s})$ and $\mathbf{Z}_{2}=f_{\theta}(\mathbf{X}_{t},\mathbf{A}_{t})$ from the shared graph encoder, we project them into spectral domains using graph Laplacian eigenvectors, yielding descriptors $\hat{\mathbf{F}}_{1}$ and $\hat{\mathbf{F}}_{2}$. The functional maps $\mathbf{C}_{12}\in\mathbb{R}^{k\times k}$ and $\mathbf{C}_{21}\in\mathbb{R}^{k\times k}$ align these spectral features via:
$\mathcal{L}_{\text{FM}}^{12}=\alpha\left\|\mathbf{C}_{12}\hat{\mathbf{F}}_{1}-\hat{\mathbf{F}}_{2}\right\|_{F}^{2}+\beta\left\|\Lambda_{2}\mathbf{C}_{12}-\mathbf{C}_{12}\Lambda_{1}\right\|_{F}^{2}$ (5)
$\mathcal{L}_{\text{FM}}^{21}=\alpha\left\|\mathbf{C}_{21}\hat{\mathbf{F}}_{2}-\hat{\mathbf{F}}_{1}\right\|_{F}^{2}+\beta\left\|\Lambda_{1}\mathbf{C}_{21}-\mathbf{C}_{21}\Lambda_{2}\right\|_{F}^{2}$ (6)
We incorporate these objectives as differentiable loss terms, treating $\mathbf{C}_{12}$ and $\mathbf{C}_{21}$ as trainable parameters that are jointly optimized end-to-end via backpropagation.
These losses are combined with the standard GAE reconstruction loss, which minimizes binary cross-entropy between the observed adjacency matrix $\mathbf{A}$ and its reconstruction $\hat{\mathbf{A}}$,
$\mathcal{L}_{\text{rec}}=\mathrm{BCE}(\mathbf{A}_{s},\hat{\mathbf{A}}_{s})+\mathrm{BCE}(\mathbf{A}_{t},\hat{\mathbf{A}}_{t}),$ (7)
where $\mathrm{BCE}(\cdot,\cdot)$ denotes the element-wise binary cross-entropy loss. The overall training objective combines the training loss and regularization terms in a weighted sum:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{FM}}\left(\mathcal{L}_{\text{FM}}^{12}+\mathcal{L}_{\text{FM}}^{21}\right)+\lambda_{\text{bij}}\,\mathcal{L}_{\text{bij}}+\lambda_{\text{orth}}\,\mathcal{L}_{\text{otrh}}$ (8)
The entire architecture is trained end-to-end via gradient descent, ensuring that functional maps and embeddings co-evolve to produce structure-aware cross-graph correspondences.
3.6 Node alignment
Given learned embeddings, we compute the cosine similarity matrix $\mathbf{S}=\hat{\mathbf{Z}}_{s}\hat{\mathbf{Z}}_{t}^{\top}$ between $\ell_{2}$-normalized node embeddings $\hat{\mathbf{Z}}_{s},\hat{\mathbf{Z}}_{t}\in\mathbb{R}^{N\times d}$ from source and target graphs. Node correspondences are predicted using greedy matching, iteratively selecting the highest similarity unmatched pairs until complete one-to-one alignment is achieved.
4 Experiments
In this section, we aim to address the following research questions: (1) robustness: is GADL more robust to feature and structural inconsistencies compared to existing state-of-the-art graph alignment methods? (2) effectiveness: does GADL outperform state-of-the-art methods on real-world graph alignment tasks? (3) generalization: How effectively can GADL generalize to vision–language alignment task? (4) ablation analysis: what is the contribution of each component in GADL to the overall alignment performance? (5) encoder evaluation: how does the proposed dual-pass GCN encoder improve node embeddings compared to standard GNN encoders? (6) hyperparameter sensitivity: How sensitive is GADL to variations in hyperparameter settings?
A comprehensive description of the experimental setup—including benchmarks, baselines, evaluation metrics, and implementation details—is provided in the Appendix C.
4.1 Robustness: evaluation on semi-synthetic benchmarks
To evaluate the robustness of the proposed GADL model under varying degrees of structural inconsistencies, we conduct experiments on six semi-synthetic graph alignment benchmark datasets: Celegans, Arena, Douban, Cora, DBLP, and Coauthor CS.
We follow the experimental setting introduced in the (He et al., 2024) for generating inconsistent graph pairs. Given a source graph $\mathcal{G}_{s}$ with adjacency matrix $\mathbf{A}$, we construct 10 perturbed and permuted target graphs using the transformation $\hat{\mathbf{A}}=\mathbf{P}(\mathbf{A}+\mathbf{M})\mathbf{P}^{\top}$, where $\mathbf{M}\in\{-1,0,1\}^{N\times N}$ introduces edge-level perturbations, and $\mathbf{P}$ is a random permutation matrix. The perturbation level is controlled by a parameter $p\in\{0,1\%,5\%\}$, representing the fraction of edges modified: $p|\mathcal{E}|$. We adopt seven structural node features from (Berlingerio et al., 2013): node degree, clustering coefficient, average degree of neighbors, average clustering coefficient of neighbors, number of edges in ego-network, number of outgoing edges of ego-network, and number of neighbors of ego-network. These descriptors provide compact, structure-aware representations for robust evaluation under varying structural inconsistencies.
Table 1 presents a comprehensive comparison of our proposed method, GADL, against six state-of-the-art network alignment methods: Netsimile (Berlingerio et al., 2013), Final (Zhang and Tong, 2016), GAlign (Trung et al., 2020), WAlign (Gao et al., 2021b), GAE (Kipf and Welling, 2016), T-GAE(He et al., 2024), and SLOTAlign (Tang et al., 2023a). Results show mean matching accuracy and standard deviation across 10 randomly generated target graphs under structural perturbations of 0%, 1%, and 5%. Results for Final, WAlign, and GAE are from He et al. (2024), while GAlign and SLOTAlign results are reproduced from their original implementations. Entries marked "–" indicate methods that failed to complete within reasonable time limits due to scalability constraints.
Table 1: Robustness evaluation under different structural inconsistency levels (%).
Dataset Perturb. NetSimile Final GAlign WAlign GAE T-GAE SLOTAlign GADL
0% 72.7 ± 0.9 92.2 ± 1.2 81.67 ± 0.7 88.4 ± 1.6 86.3 ± 1.3 91.0 ± 1.1 91.12 ± 0.2 92.82 ± 0.9
Celegans 1% 66.3 ± 3.8 33.2 ± 7.8 66.23 ± 0.8 80.7 ± 3.0 33.2 ± 8.4 86.5 ± 1.1 85.25 ± 0.6 88.07 ± 0.7
5% 41.1 ± 13.0 10.4 ± 2.7 49.22 ± 1.6 42.4 ± 21.1 6.5 ± 2.4 69.2 ± 2.1 70.05 ± 0.4 71.74 ± 0.2
0% 94.7 ± 0.3 97.5 ± 0.3 93.02 ± 0.4 97.4 ± 0.5 97.6 ± 0.4 97.8 ± 0.4 96.22 ± 0.5 98.27 ± 0.3
Arena 1% 87.8 ± 1.0 32.5 ± 5.9 87.46 ± 0.6 90.0 ± 3.1 30.1 ± 17.6 96.0 ± 1.0 95.24 ± 0.4 96.86 ± 0.4
5% 52.3 ± 5.3 7.2 ± 2.6 64.96 ± 1.2 30.4 ± 17.5 1.4 ± 1.4 78.6 ± 2.5 78.5 ± 0.6 80.69 ± 0.4
0% 46.4 ± 0.4 89.9 ± 0.3 56.50 ± 1.4 90.0 ± 0.4 89.5 ± 0.4 90.1 ± 0.3 88.17 ± 0.3 90.71 ± 0.4
Douban 1% 40.0 ± 1.2 27.8 ± 5.7 51.40 ± 0.5 77.2 ± 4.8 38.3 ± 16.4 87.3 ± 0.4 85.83 ± 1.2 87.94 ± 0.1
5% 20.7 ± 4.6 7.8 ± 3.0 29.97 ± 2.2 36.6 ± 13.4 0.6 ± 0.3 70.2 ± 2.5 67.42 ± 0.5 69.62 ± 0.1
0% 73.7 ± 0.4 87.5 ± 0.7 74.15 ± 0.7 87.2 ± 0.4 87.1 ± 0.8 87.5 ± 0.4 87.74 ± 0.6 88.20 ± 0.2
Cora 1% 66.4 ± 1.6 30.0 ± 3.3 68.53 ± 0.4 80.1 ± 1.2 57.9 ± 5.3 85.1 ± 0.5 84.66 ± 0.1 85.30 ± 0.3
5% 41.2 ± 3.3 6.7 ± 2.8 45.67 ± 0.8 33.4 ± 7.3 9.6 ± 2.7 67.7 ± 1.3 67.8 ± 0.3 68.22 ± 0.2
0% 63.7 ± 0.2 85.6 ± 0.2 66.43 ± 0.6 85.6 ± 0.2 85.2 ± 0.3 85.6 ± 0.2 85.82 ± 0.0
DBLP 1% 55.1 ± 1.7 15.2 ± 3.3 59.00 ± 0.5 73.1 ± 1.6 19.4 ± 0.6 83.3 ± 0.4 82.77 ± 0.3
5% 19.5 ± 4.8 2.7 ± 0.9 38.84 ± 0.2 15.9 ± 8.3 1.4 ± 0.2 60.8 ± 1.9 62.49 ± 0.3
0% 90.9 ± 0.1 97.6 ± 0.1 92.18 ± 1.5 97.5 ± 0.2 97.6 ± 0.3 97.6 ± 0.1 97.76 ± 0.1
Coauthor CS 1% 75.2 ± 2.2 13.3 ± 5.0 81.15 ± 0.7 75.2 ± 5.4 49.5 ± 7.8 93.2 ± 0.8 93.41 ± 0.6
5% 26.3 ± 6.0 2.0 ± 0.4 30.41 ± 0.1 11.3 ± 7.5 0.6 ± 0.1 66.0 ± 1.4 68.54 ± 1.2
The experimental results yield several key observations: (1) GADL consistently ranks among the top performers across all datasets, demonstrating robust performance under varying structural inconsistencies. While baseline methods often exhibit sharp performance degradation under structural noise, GADL maintains high accuracy even with 5% edge perturbations. (2) Embedding-based GA methods (e.g., GAE, T-GAE, GAlign, GADL) generally outperform optimal-transport-based methods (e.g., Final). SLOTAlign, which combines both learning-based and optimization techniques, achieves competitive but suboptimal results compared to pure learning-based models.
4.2 Effectiveness: evaluation on real-world benchmarks
We evaluate the effectiveness of the proposed GADL method on two real-world noisy graph datasets with partial node alignment: Douban Online-Offline and ACM-DBLP. These benchmarks reflect realistic and challenging scenarios, involving distinct graphs with partially aligned nodes. We evaluate the performance of GADL and baseline methods using Hit@$k$, which measures the proportion of ground-truth aligned nodes ranked among the top-$k$ predictions. The results are reported in Table 2.
Table 2: Performance of graph alignment methods on real-world benchmarks.
Method ACM-DBLP Douban Online-Offline
Hit@1 Hit@5 Hit@10 Hit@50 Hit@1 Hit@5 Hit@10 Hit@50
NetSimile 2.59 8.32 12.09 26.42 1.07 2.77 4.74 15.03
GAE 8.10 22.50 30.10 45.10 3.30 9.20 14.10 32.10
GAlign 73.26 91.24 95.09 98.37 41.32 62.43 71.37 87.65
WAlign 62.02 81.96 87.31 93.89 36.40 53.94 67.08 85.33
T-GAE 73.89 91.73 95.33 98.22 36.94 60.64 69.77 88.62
SLOTAlign 66.04 84.06 87.95 94.65 51.43 53.43 77.73 90.23
GADL 88.63 94.76 96.16 98.41 53.31 73.61 80.67 94.18
The results reveal the following key insights: (1) GADL consistently achieves the highest alignment accuracy across both real-world datasets, outperforming all baselines. In particular, it surpasses the second-best models—T-GAE on ACM-DBLP and SLOTAlign on Douban—by substantial margins, demonstrating its robustness in unsupervised real-world graph alignment scenarios. (2) Compared to T-GAE, which employs a GAE framework with a GIN encoder but lacks latent-space communication, our GADL model demonstrates superior performance through its dual-pass GCN encoder architecture integrated with latent-space communication. This design enables GADL to generate superior node embeddings, improving performance on unsupervised graph alignment tasks. (3) Beyond GADL, other learning-based methods (e.g., T-GAE) also outperform optimal-transport approaches (e.g., SLOTAlign) on larger benchmarks (e.g., ACM-DBLP), demonstrating superior robustness of learning-based methods to structural variations in large graphs that violate optimal transport assumptions.
4.3 Generalization: evaluation on vision–language benchmarks
Latent space alignment is a special case of graph alignment, relying only on embeddings without explicit structure. To highlight this generality, we further evaluate our method on vision-language alignment benchmarks, where the task involves aligning latent representations from diverse pretrained vision and language models.
We evaluate latent space alignment across multiple benchmarks using representations from diverse pretrained vision and language models. Full experimental details are provided in Appendix C.4.
Table 3 summarizes the vision-language alignment accuracies on four datasets using three pretrained vision models (CLIP (Ramesh et al., 2022), DeiT (Touvron et al., 2021), and DINOv2 (Oquab et al., 2023)) and two pretrained language models from SentenceTransformers library (Reimers and Gurevych, 2019) (all-mpnet-base-v2 and all-roberta-large-v1). Comprehensive results for all vision and language model combinations are provided in the Appendix F. Results on CIFAR-100 and ImageNet-100 are reproduced using official implementations.
Table 3: Vision-language alignment across four datasets using three pretrained vision models (CLIP, DeiT, and DINOv2) and two pretrained language models: Lan. model 1 (all-mpnet-base-v2) and Lan. model 2 (all-roberta-large-v1).
Method CIFAR-10 CINIC-10 CIFAR-100 ImageNet-100
Lan. model 1 Lan. model 2 Lan. model 1 Lan. model 2 Lan. model 1 Lan. model 2 Lan. model 1 Lan. model 2
CLIP - (ViT-L/14@336)
LocalCKA 25.0 ± 10.5 17.0 ± 15.9 30.0 ± 0.0 4.0 ± 5.0 24.00 ± 1.41 13.67 ± 0.47 8.00 ± 1.41 8.33 ± 0.47
OT 0.0 ± 0.0 10.0 ± 0.0 49.5 ± 2.2 2.0 ± 4.1 1.00 ± 0.00 1.67 ± 0.47 1.33 ± 0.47 1.00 ± 0.00
FAQ 12.0 ± 10.1 0.5 ± 2.2 30.5 ± 2.2 0.0 ± 0.0 2.33 ± 1.25 2.67 ± 1.70 4.33 ± 1.70 2.33 ± 1.25
MPOpt 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 1.67 ± 1.25 2.67 ± 1.25 4.67 ± 2.05 2.67 ± 0.47
Gurobi 20.5 ± 6.0 47.0 ± 7.3 50.0 ± 0.0 80.0 ± 0.0 2.11 ± 1.29 3.44 ± 1.27 3.22 ± 2.35 4.50 ± 1.50
Hahn-Grant 25.0 ± 10.5 47.0± 7.3 50.0 ± 0.0 80.0 ± 0.0 2.33 ± 1.25 3.00 ± 2.16 4.93 ± 2.05 4.67 ± 1.70
GADL 76.7 ± 4.7 73.3 ± 4.7 76.7 ± 4.7 80.0 ± 0.0 79.7 ± 2.0 81.00 ± 1.4 41.3 ± 8.2 45.3 ± 14.2
DeiT - (DeiT-B/16d@384)
LocalCKA 24.0 ± 9.9 20.0 ± 5.6 68.0 ± 8.9 0.0 ± 0.0 10.33 ± 0.94 23.33 ± 0.47 8.33 ± 1.70 9.33 ± 0.47
OT 12.0 ± 4.1 10.0 ± 0.0 20.0 ± 0.0 0.0 ± 0.0 2.33 ± 0.94 1.67 ± 0.47 2.00 ± 0.00 0.67 ± 0.47
FAQ 40.0 ± 15.2 22.5 ± 9.7 55.5 ± 5.1 0.0 ± 0.0 4.33 ± 0.47 1.33 ± 1.25 3.67 ± 0.47 3.33 ± 0.94
MPOpt 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.33 ± 0.47 0.67 ± 0.94 2.67 ± 2.36 1.00 ± 0.82
Gurobi 28.5 ± 3.7 59.0 ± 3.1 10.0 ± 0.0 40.0 ± 0.0 3.67 ± 2.49 3.11 ± 1.91 3.56 ± 1.57 3.00 ± 1.00
Hahn-Grant 28.5 ± 3.7 59.0 ± 3.1 10.0 ± 0.0 40.0 ± 0.0 1.33 ± 0.47 5.33 ± 1.25 1.67 ± 2.36 1.33 ± 1.25
GADL 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 47.3 ± 8.9 42.7 ± 6.3 67.3 ± 4.5 65.7 ± 0.9
DINOv2 - (ViT-G/14)
LocalCKA 37.5 ± 28.8 18.5 ± 29.2 52.5 ± 31.1 57.0 ± 13.4 4.00 ± 0.82 4.67 ± 0.94 5.33 ± 0.47 6.00 ± 0.82
OT 30.0 ± 13.8 33.5 ± 19.8 77.5 ± 6.4 15.5 ± 7.6 1.00 ± 0.00 1.00 ± 0.00 1.00 ± 0.00 0.33 ± 0.47
FAQ 37.5 ± 21.2 38.0 ± 29.8 31.0 ± 4.5 29.5 ± 2.2 4.33 ± 0.94 3.33 ± 1.25 3.00 ± 0.82 4.33 ± 2.05
MPOpt 73.5 ± 17.9 94.0 ± 18.5 79.0 ± 3.1 47.0 ± 46.0 1.33 ± 1.25 0.33 ± 0.47 4.00 ± 0.82 0.67 ± 0.47
Gurobi 69.5 ± 24.2 100.0 ± 0.0 79.0 ± 3.1 100.0 ± 0.0 2.56 ± 1.64 1.78 ± 1.31 1.50 ± 0.50 2.50 ± 0.50
Hahn-Grant 69.5 ± 24.2 100.0 ± 0.0 79.0 ± 3.1 100.0 ± 0.0 4.00 ± 0.82 2.00 ± 1.41 6.33 ± 0.47 1.22 ± 0.92
GADL 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 67.7 ± 1.7 58.3 ± 0.4 44.3 ± 8.3 49.3 ± 0.4
The experimental results highlight several key insights. (1) GADL consistently outperforms all baselines across datasets and pretrained models, demonstrating that moving beyond optimization- and optimal transport–based frameworks provides substantial benefits. (2) Most baselines achieve near-chance accuracies ($\leq 10\%$) with occasional inconsistent successes, and even sophisticated solvers like Gurobi or Hahn-Grant fail in certain settings. This inconsistency intensifies with dataset complexity, and the performance gap with GADL becomes pronounced on challenging benchmarks (CIFAR-100, ImageNet-100), underscoring the limitations of treating vision-language alignment as pure assignment optimization. (3) The choice of pretrained model critically impacts cross-modal alignment performance. While DINOv2 and DeiT achieve highest accuracies on smaller datasets, CLIP consistently outperforms on larger, more complex benchmarks. This highlights how pretrained model selection significantly influences visual-language alignment effectiveness.
4.4 Ablation analysis
To better understand the contribution of each component in the proposed GADL framework, we conduct an ablation study evaluating variants with specific modules removed or modified. We compare GADL against: (1) GADL w/o dual-pass encoder: replaces the dual-pass GCN with a standard single-pass GCN while retaining latent-space communication; (2) GADL w/o bijectivity regularization: removes the bijectivity regularization term while keeping the dual-pass encoder; (3) GADL w/o orthogonality regularization: omits orthogonality regularization while maintaining all other components.
Table 4: Performance of GADL and its variants on real-world graph alignment benchmarks.
Method ACM-DBLP Douban Online-Offline
Hit@1 Hit@5 Hit@10 Hit@50 Hit@1 Hit@5 Hit@10 Hit@50
GADL w/o dual-pass encoder 81.68 92.22 95.24 97.88 43.38 62.96 71.1 88.55
GADL w/o bijectivity regularization 88.47 94.6 96.06 98.37 52.68 72.89 80.14 94.78
GADL w/o orthogonality regularization 88.51 94.48 96.06 98.35 51.96 72.8 79.51 94.72
GADL 88.63 94.76 96.16 98.41 53.31 73.61 80.67 94.18
The ablation results are summarized in Table 4. They highlight the individual contribution of each component to the overall alignment performance. Replacing the dual-pass GCN with a standard encoder causes substantial accuracy drops. Furthermore, the removal of the regularization terms notably degrades performance, confirming their importance in achieving robust and reliable node alignment.
The results also show that the dual-pass encoder provides greater improvement on Douban than ACM-DBLP, reflecting the impact of initial node features on encoder effectiveness. Essentially, the Douban dataset contains sparse, high-dimensional node features with many zero entries, causing standard GCN embeddings to become overly smooth and less discriminative. Consequently, the dual-pass filters lead to a significant improvement in matching accuracy. In contrast, the ACM-DBLP features are denser and more informative, so the standard GCN already generates sufficiently distinctive embeddings, with the high-pass component providing only moderate improvement.
Further analysis and experiments can be found in the supplementary material (Appendices D and E).
5 Conclusion
We present GADL, a novel framework for unsupervised graph alignment that combines dual-pass encoding with geometry-aware latent space communication. Comprehensive experiments demonstrate consistent performance gains across diverse benchmarks, with successful application to vision-language tasks validating the framework’s broader utility. While promising, our approach incurs modest computational overhead from dual-pass encoding compared to standard GCN and requires careful hyperparameter tuning. Future work will focus on adaptive spectral filtering and efficient embedding strategies, with potential extensions to molecular networks, social graphs, and multi-modal alignment tasks.
Appendices
Appendix A Related work
Extensive research has addressed the graph alignment problem, with existing methods broadly categorized into three families based on their alignment strategies: optimization-based, optimal transport-based, and embedding-based approaches. These methods also differ in terms of the level of supervision required—ranging from unsupervised techniques to semi-supervised methods (which rely on partially paired nodes), and fully supervised approaches.
Traditional graph alignment methods formulate the problem as an optimization task, typically as a Quadratic Assignment Problem (QAP), seeking node permutations that minimize discrepancies between source and target adjacency matrices. IsoRank [Singh et al., 2008] represents a seminal approach, employing a PageRank-inspired algorithm to compute node similarity matrices based on neighbor similarity for unsupervised alignment. BigAlign [Koutra et al., 2013] extends this framework by incorporating both structural and attribute information to enhance alignment accuracy. FINAL [Zhang and Tong, 2016] addresses scalability through matrix factorization, combining global structural consistency with partial anchor constraints.
These optimization-based approaches often struggle with scalability due to the NP-hard nature of QAP, though approximation strategies and relaxations can make them tractable on medium-sized networks. While primarily unsupervised, they can accommodate semi-supervised settings by incorporating known anchor pairs as hard or soft constraint
Optimal transport-based methods model each graph as a probability distribution over its nodes and seek a transport plan—i.e., a soft correspondence—that minimizes a divergence such as the Wasserstein or Gromov-Wasserstein distance between the distributions. This framework offers a principled approach to graph alignment by optimizing the transport cost between node distributions. Unlike the hard alignments produced by QAP-based methods, optimal transport typically yields soft alignment matrices, allowing for uncertainty and partial correspondences.
A notable early contribution in this category is WAlign [Gao et al., 2021b], which jointly learns node embeddings and alignments by minimizing Wasserstein distance between graphs in a shared embedding space using a lightweight GCN and Wasserstein distance discriminator. Building on this direction, FGW [Tang et al., 2023b] employs Fused Gromov-Wasserstein distance to jointly align structural and attribute information through a coarse-to-fine matching scheme. PARROT [Zeng et al., 2023] extends this idea by running Random Walk with Restart (RWR) on both individual graphs and their Cartesian product, capturing more nuanced structural correspondence. GALOPA [Wang et al., 2023] integrates a GNN encoder with a self-supervised OT loss, jointly learning features and transport plans. To improve scalability, Wasserstein Wormhole [Haviv et al., 2024] introduces a transformer-based autoencoder that maps distributions into a latent space where Euclidean distances approximate Wasserstein distances, enabling efficient, linear-time graph comparisons.
These methods are typically unsupervised and particularly effective for noisy or incomplete graphs due to their probabilistic formulation and global alignment perspective.
Embedding-based methods learn vector representations for nodes in each graph and align them based on embedding similarity. This approach typically involves generating node embeddings—either independently or jointly—followed by alignment through nearest-neighbor search or learned mapping functions.
NetSimile [Berlingerio et al., 2013] represents an early embedding-based approach that uses handcrafted structural features (degree, clustering coefficient) to represent nodes and aligns graphs through direct feature vector comparison using similarity measures. GAlign [Trung et al., 2020] adopts an unsupervised approach where both graphs are independently encoded using a shared Graph Convolutional Network, with node embeddings aligned by minimizing distributional discrepancies such as Wasserstein distance between embedding spaces. NeXtAlign [Zhang et al., 2021] enhances representation learning through a cross-graph attention mechanism that enables nodes in one graph to attend to features in the other. This produces alignment-aware embeddings and improves performance in semi-supervised settings with known anchor node pairs. REGAL [Heimann et al., 2018] generates compact node embeddings by extracting structural features like node degree and local neighborhoods, then aligns nodes across graphs by matching their embeddings based on distance, enabling efficient and scalable graph alignment. GINA [Wang et al., 2022b] addresses hierarchical alignment by projecting node embeddings from Euclidean to hyperbolic space, learning linear transformations between geometries using anchor nodes to better capture scale-free and hierarchical structures in social and biological networks.
A foundational approach to embedding-based graph learning is the Graph Autoencoder (GAE) and its probabilistic extension, the Variational Graph Autoencoder (VGAE) [Kipf and Welling, 2016]. These models use a GCN encoder to generate latent node embeddings, which are then used to reconstruct the adjacency matrix via an inner product decoder. Although originally designed for link prediction, GAEs have become a common backbone for alignment tasks due to their ability to capture global graph structure in an unsupervised manner. Expanding on this foundation, T-GAE [He et al., 2024] addresses scalability through a transferable graph autoencoder trained on small graph families that generalizes to large, unseen networks without fine-tuning. This design enables strong alignment performance while significantly reducing training time and computational overhead. However, typical embedding-based methods often become unstable when graphs differ significantly in structure. SLOTAlign [Tang et al., 2023a] is developed to tackle the structure and feature inconsistencies commonly found in these embedding-based graph alignment methods. It formulates alignment as an optimal transport problem on learned intra-graph similarity matrices, combining optimal transport with embedding-based approaches.
Appendix B Proof of Theorem 1
Proof.
Preliminary Definitions: Let GCN filter $\tilde{\mathbf{A}}_{\text{GCN},\text{sym}}=\mathbf{I}-\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}(\mathbf{I}-\tilde{\mathbf{\Lambda}})\mathbf{U}^{\top}$, where $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{n}]$ is the eigenbasis, $\boldsymbol{\hat{\Lambda}}=\text{diag}(\hat{\lambda}_{1},\ldots,\hat{\lambda}_{n})$ is the diagonal matrix of eigenvalues, and each eigenvalue satisfies $\hat{\lambda}_{i}\in[0,2]$. The spectral representation of node features is: $\mathbf{X}=\sum_{k=1}^{n}\hat{\mathbf{X}}_{k}\mathbf{u}_{k}$, where $\hat{\mathbf{X}}_{k}=\mathbf{u}_{k}^{\top}\mathbf{X}$.
  • The low-pass component captures the smooth, global structure of the graph. It aggregates information from neighbors, producing embeddings: $\mathbf{Z}_{\text{low}}=\sum_{k=1}^{n}p_{\text{low}}(\hat{\lambda}_{k})\,\hat{\mathbf{X}}_{k}\,\mathbf{u}_{k}$, where $p_{\text{low}}(\hat{\lambda}_{k})=1-\frac{1}{2}\hat{\lambda}_{k}$. This captures smoothed signals over the graph—node embeddings are averages of their neighbors.
  • The high-pass component captures complementary, local variations and finer structural details, given by: $\mathbf{Z}_{\text{high}}=\sum_{k=1}^{n}p_{\text{high}}(\hat{\lambda}_{k})\,\hat{\mathbf{X}}_{k}\,\mathbf{u}_{k}$, where $p_{\text{high}}(\hat{\lambda}_{k})=\frac{1}{2}\hat{\lambda}_{k}$.
Claim 1 (Neighborhood preservation). The dual-pass embedding $\mathbf{z}_{i}$ preserves neighborhood similarity as effectively as the low-pass embedding $\mathbf{z}_{i}^{\text{low}}$.
Proof of Claim 1. The key insight underlying local consistency is that neighborhood similarity is primarily encoded in low-frequency spectral components, which capture smooth variations across connected nodes.
For the dual-pass embeddings of nodes $i$ and $j$, the cosine similarity can be written as:
$\langle\mathbf{z}_{i},\mathbf{z}_{j}\rangle=\langle\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{low}}\rangle+\langle\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{high}}\rangle+\langle\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{high}}\rangle+\langle\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{low}}\rangle.$
The filters are designed to be spectrally complementary,
$p_{\text{low}}(\hat{\lambda}_{k})+p_{\text{high}}(\hat{\lambda}_{k})=(1-\frac{1}{2}\hat{\lambda}_{k})+\frac{1}{2}\hat{\lambda}_{k}=1,\quad\forall\hat{\lambda}_{k}\in[0,2].$
This spectral separation implies that $\mathbb{E}\big{[}\langle\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{i}^{\text{high}}\rangle\big{]}\approx 0$, meaning that the low-pass and high-pass components are approximately orthogonal in expectation.
Therefore, the low-pass and high-pass components are approximately orthogonal, and the high-pass component adds complementary information without significantly affecting the neighborhood-preserving properties encoded by the low-pass component. Consequently, the dual-pass embedding preserves neighborhood similarity as effectively as the low-pass component alone.
Claim 2 (Enhanced discriminability for node correspondence). For node correspondence tasks, the dual-pass embedding $\mathbf{z}_{i}$ provides superior discriminability compared to either $\mathbf{z}_{i}^{\text{low}}$ or $\mathbf{z}_{i}^{\text{high}}$ alone. That is, false correspondences are less likely under similarity computed via $\mathbf{z}_{i}$.
Proof of Claim 2. Discriminability is measured by the separation margin between the distributions of similarities for corresponding pairs $\mathcal{C}=\{(i,j):i\in\mathcal{V}_{1},\;j\in\mathcal{V}_{2},\;i\leftrightarrow j\}$ and non-corresponding pairs $\mathcal{N}=\{(i,j):i\in\mathcal{V}_{1},\;j\in\mathcal{V}_{2},\;i\not\leftrightarrow j\}$.
As we mentioned, the dual-pass filter design ensures perfect spectral complementarity: the low-pass and high-pass filters have anti-correlated frequency responses, ensuring that their contributions are nearly independent. Consequently, their mutual information is bounded, $I(\mathbf{z}_{i}^{\text{low}};\mathbf{z}_{i}^{\text{high}})\leq\epsilon$, for small $\epsilon>0$, reflecting the opposing spectral emphasis.
For any $\ell_{2}$-induced metric, the squared distance between dual-pass embeddings decomposes as:
$d(\mathbf{z}_{i},\mathbf{z}_{j})^{2}=d(\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{low}})^{2}+d(\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{high}})^{2}+2\langle\mathbf{z}_{i}^{\text{low}}-\mathbf{z}_{j}^{\text{low}},\mathbf{z}_{i}^{\text{high}}-\mathbf{z}_{j}^{\text{high}}\rangle.$
Under the approximate orthogonality condition, the cross-term is negligible, yielding:
$d(\mathbf{z}_{i},\mathbf{z}_{j})^{2}\approx d(\mathbf{z}_{i}^{\text{low}},\mathbf{z}_{j}^{\text{low}})^{2}+d(\mathbf{z}_{i}^{\text{high}},\mathbf{z}_{j}^{\text{high}})^{2}.$
The separation margin is thus defined as $\Delta=\min_{(i,j)\in\mathcal{C}}\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})-\max_{(i,k)\in\mathcal{N}}\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})$.
The critical observation is that the two components offer complementary discriminative power:
  • Case A (Low-pass insufficient): When graphs share similar global structure but differ in local details, $\mathbf{z}_{i}^{\text{low}}$ may yield high similarity for non-corresponding pairs. However, $\mathbf{z}_{i}^{\text{high}}$ captures local differences, reducing false positives.
  • Case B (High-pass insufficient): When local structures are noisy or similar, $\mathbf{z}_{i}^{\text{high}}$ may be unreliable. However, $\mathbf{z}_{i}^{\text{low}}$ provides stable global discrimination based on community structure and smooth attributes.
Together, these effects yield an additive improvement in discriminability. Formally, the dual-pass margin satisfies
$\Delta_{\text{dual}}\geq\max(\Delta_{\text{low}},\Delta_{\text{high}})+\gamma,$
where $\gamma>0$ represents the additional discriminative contribution from orthogonal spectral information. This establishes that the dual-pass embedding provides superior node correspondence discrimination.
Appendix C Experimental setup
This section describes the benchmarks, performance metrics, and experimental settings used for graph alignment evaluation.
C.1 Benchmarks
Table 5 summarizes statistics for all experimental datasets. It includes six semi-synthetic graph alignment benchmarks consisting of graphs with varying sizes and properties to comprehensively evaluate the robustness of our approach. Additionally, two real-world graph alignment datasets with partial ground-truth node correspondences are included to assess overall performance. The datasets are as follows:
  • Celegans: This dataset models the protein-protein interaction network of Caenorhabditis elegans. Each node represents a protein, and edges indicate physical or functional interactions between proteins, making it useful for biological network analysis and alignment tasks involving molecular networks [Kunegis, 2013].
  • Arenas: A communication network derived from email exchanges at the University Rovira i Virgili. Nodes correspond to individual users, and edges represent the presence of at least one email sent between them. It serves as a social interaction graph with temporal and communication patterns [Leskovec and Krevl, 2014].
  • Douban: A social network from the Chinese movie review platform Douban, where nodes represent users, and edges capture friend or contact relationships. This dataset is commonly used to study social dynamics and network alignment in social media contexts [Zhang and Tong, 2016].
  • Cora: A citation network of scientific papers where nodes are publications, and edges denote citation relationships. Cora is a benchmark dataset for graph mining and node classification, providing a structured academic citation graph ideal for evaluating graph-based learning models [Sen et al., 2008].
  • DBLP: An extensive citation network aggregated from DBLP, Association for Computing Machinery (ACM), Microsoft Academic Graph (MAG), and other scholarly databases. It includes publication and citation information, widely used for testing graph alignment, clustering, and knowledge discovery tasks in academic networks [Pan et al., 2016].
  • CoauthorCs: A co-authorship network in computer science that represents collaborations between authors. Nodes correspond to researchers, and edges indicate joint publications. It is often used to study community structure and author disambiguation in bibliographic databases [Sinha et al., 2015].
  • ACM-DBLP: This dataset contains two co-authorship graphs from the ACM and DBLP databases. Nodes represent authors, and edges indicate co-authorship. Although collected independently, both graphs share overlapping authors, with 6,325 ground-truth alignments. Node features capture publication distributions across research venues. The dataset poses a challenging alignment task due to structural and feature discrepancies between the two graphs [Zhang and Tong, 2018].
  • Douban Online-Offline: This dataset comprises two social graphs from the Douban platform—one based on online interactions and the other on offline event co-attendance. Both graphs share a subset of users, with 1,118 aligned nodes. Node features reflect user location distributions. The dataset is designed to evaluate alignment across heterogeneous and partially overlapping social networks [Zhang and Tong, 2016].
Table 5: Overview of datasets and their key properties
Dataset #Nodes #Edges #Aligned Nodes #Node Features Description
Celegans 453 2025 453 7 Interactome
Arenas 1133 5451 1133 Email Communication
Douban 3906 7215 3906 Social Network
Cora 2708 5278 2708 Citation Network
DBLP 17716 52867 17716 Citation Network
CoauthorCs 18333 81894 18333 Coauthor
ACM-DBLP ACM 9872 39561 6325 17 Coauthor Network
DBLP 9916 44808 Coauthor Network
Douban Online 3906 16328 1118 538 Social Network
Offline 1118 3022 Social Network
C.2 Performance metrics
To assess graph alignment performance, we adopt two widely used metrics: alignment accuracy (Acc) and Hit@$k$. Alignment Accuracy (Acc) measures the proportion of correctly predicted node correspondences among all ground-truth aligned pairs, providing a direct measure of overall matching performance. Hit@$k$ evaluates whether the true corresponding node from the source graph appears within the top-$k$ predicted candidates for each node in the target graph. This metric reflects the ability of model to rank correct matches highly and is particularly useful for top-$k$ retrieval scenarios.
Both metrics are computed using all available ground-truth node pairs, with higher values indicating better alignment quality.
C.3 Implementation details
As shown in Figure 2, our proposed architecture consists of two GAEs, each equipped with a dual-pass GCN encoder. The number of GCN layers is selected based on empirical performance for each benchmark: 2 layers for ACM-DBLP and the semi-synthetic datasets, 3 for DBLP, 5 for Douban, and 6 for Douban Online-Offline.
For computing functional maps, we use the first 300 eigenvectors of the graph Laplacian across all datasets. The hidden dimensions of the encoder are set to 1024 for ACM-DBLP, 256 for Douban Online-Offline, and 16 for all semi-synthetic benchmarks. For all baseline methods, we report results from the original papers when available. For cases where published results were not available on specific benchmarks (notably for GAlign and SLOTAlign in Section 4.1), we re-ran the official code with default configurations to ensure a fair comparison. The model is trained end-to-end using the Adam optimizer with a learning rate of 1e-3 and a weight decay of 5e-4, based on the loss function in Equation 8.
All experiments are implemented using PyTorch 2.1.2 and PyTorch Geometric 2.5.0. Most benchmarks are run on servers equipped with NVIDIA A100 GPUs (CUDA 12.2), each providing 40 GB of memory. For large-scale datasets such as DBLP and Coauthor CS, we use NVIDIA H100 GPUs (CUDA 12.6) with 95 GB of memory, enabling efficient training and evaluation on high-complexity graphs.
C.4 Details of vision–language experiment
C.4.1 Setup
We evaluate the vision-language alignment task on a range of benchmarks, including CIFAR-10 [Krizhevsky et al., 2009], CINIC-10 [Darlow et al., 2018], CIFAR-100 [Krizhevsky et al., 2009], and ImageNet-100 [Russakovsky et al., 2015], using representations extracted from diverse pretrained vision and language models. For each vision model, class-level representations are derived by averaging image-level embeddings within each class. Correspondingly, language representations are obtained by averaging embeddings generated from multiple textual prompts for each class. To enable application of our graph alignment method, we build a similarity graph from these representations, where each class-level embedding is treated as a node and connected to its $k$ most similar neighbors according to cosine similarity.
C.4.2 Baselines
We compare against a set of established solvers and heuristics for the alignment problem. LocalCKA [Maniparambil et al., 2024] leverages the centered kernel alignment (CKA) metric to approximate the QAP with a linear assignment formulation, providing an efficient method for vision–language correspondence. Optimal Transport (OT) methods [Peyré et al., 2016] address the alignment by modeling embeddings as probability distributions and computing the minimal transport cost, thereby preserving geometric structure across modalities. The Fast Approximate QAP algorithm (FAQ) [Vogelstein et al., 2015] is a well-known primal heuristic that relaxes the QAP and iteratively refines the solution, yielding scalable but approximate alignments. MPOpt [Hutschenreiter et al., 2021] represents a generic mathematical programming approach, solving the alignment as a constrained optimization problem using standard formulations. Gurobi [Gurobi Optimization, LLC, 2023] is a commercial off-the-shelf solver for mixed-integer and quadratic programs, providing near-optimal results for small problem instances. Finally, the Hahn-Grant solver [Schnaus et al., 2025] is a dual ascent algorithm that produces strong lower bounds by repeatedly solving linear assignment problems.
C.4.3 Vision and language models
We adopt the set of 32 vision models used in Blind Match [Schnaus et al., 2025]. For self-supervised methods, we use DINO [Caron et al., 2021] models (RN50 and ViT-S/B with patch sizes 16 and 8) trained on ImageNet-1k and DINOv2 [Oquab et al., 2023] models (ViT-S/B/L/G with patch size 14) trained on the LVD-142M dataset, as well as fully supervised models such as DeiT variants [Touvron et al., 2021] (Tiny, Small, and Base with patch size 16, including distilled and high-resolution @384 versions) and ConvNeXt models[Liu et al., 2022b] (Base and Large, pretrained on ImageNet-1k or ImageNet-22k, with additional fine-tuned @384 variants). For vision–language pretraining, we employ CLIP [Ramesh et al., 2022] with both ResNet backbones (RN50, RN101, RN50x4, RN50x16, RN50x64) and Vision Transformer architectures (ViT-B/32, ViT-B/16, ViT-L/14, ViT-L/14@336). All experiments are conducted using official implementations and pretrained weights, ensuring consistent and reliable representation extraction for each model.
We consider four pretrained language models spanning diverse architectures and training paradigms, including the RN50x4 model from CLIP [Ramesh et al., 2022] and three models—all-MiniLM-L6-v2, all-mpnet-base-v2, and all-Roberta-large-v1—extracted from the SentenceTransformers library [Reimers and Gurevych, 2019]. All vision and language models used in our experiments are summarized in Table 6.
Table 6: Summary of vision and language models used in the experiments
Vision models
DINO [Caron et al., 2021]: RN50, ViT-S/16, ViT-S/8, ViT-B/16, ViT-B/8
DINOv2 [Oquab et al., 2023]: ViT-S/14, ViT-B/14, ViT-L/14, ViT-G/14
DeiT [Touvron et al., 2021]: DeiT-T/16, DeiT-T/16d, DeiT-S/16, DeiT-S/16d, DeiT-B/16, DeiT-B/16@384, DeiT-B/16d, DeiT-B/16d@384
ConvNeXt [Liu et al., 2022b]: CN-B-1, CN-B-22, CN-L-1, CN-L-22, CN-L-22ft@384, CN-XL-22ft@384
CLIP [Ramesh et al., 2022]: RN50, RN101, RN50x4, RN50x16, RN50x64, ViT-B/32, ViT-B/16, ViT-L/14, ViT-L/14@336
Language models
CLIP [Ramesh et al., 2022]: RN50x4
SentenceTransformers [Reimers and Gurevych, 2019]: all-MiniLM-L6-v2, all-mpnet-base-v2, all-Roberta-large-v1
Appendix D Ablation study on graph encoders
We evaluate the impact of different GNN encoder architectures on the alignment accuracy within the GADL framework using two real-world benchmark datasets. Specifically, we conduct a comparative evaluation of our proposed dual-pass GCN encoder against four GNN variants: GCN [Kipf and Welling, 2017], GIN [Xu et al., 2018a], JKGNN [Xu et al., 2018b], and TIDE [Behmanesh et al., 2023].
For these encoders, we adopt a 6-layer architecture with ReLU activation functions, following the configurations presented in their respective papers. Additionally, TIDE is applied in a single-channel setup, where the learnable parameter $t$ is shared across all channels.
Figure 3: Encoder comparison on graph alignment performance (Hit@$K$).
As shown in Figure 3, the proposed dual-pass GCN achieves consistently higher alignment accuracy across both datasets. Among the others, the GIN encoder performs best because it is designed to better capture graph structure by extending the Weisfeiler-Lehman (WL) graph isomorphism test, which helps it distinguish nodes more effectively. Since more expressive node representations reduce ambiguity in identifying correct correspondences, this enhanced expressiveness directly contributes to improved node alignment accuracy.
Appendix E Hyperparameter sensitivity
We conduct a sensitivity analysis to examine the influence of key hyperparameters on model performance. In particular, we focus on the loss weighting coefficients $\lambda_{\text{FM}}$, $\lambda_{\text{bij}}$, and $\lambda_{\text{orth}}$, which control the relative importance of the functional map loss, bijectivity constraint, and orthogonality regularization, respectively, in the overall training objective.
To isolate the effect of each hyperparameter, we vary its value over a defined range while keeping the remaining parameters fixed at their optimal values, as determined through prior validation. We evaluate model performance using the Hit@1 accuracy metric on both real-world benchmark datasets.
The results in Figure 4 demonstrate how model performance responds to variations in each hyperparameter, using the optimal values identified through tuning as a reference. The results demonstrate that our approach is stable across a wide range of settings, while also pointing out where tuning is most important for best results. These insights provide practical guidance for selecting effective hyperparameter configurations when applying the model to new benchmarks.
Figure 4: Hyperparameter sensitivity analysis on ACM-DBLP and Douban Online-Offline datasets. We report Hit@1 accuracy while varying each hyperparameter ($\lambda_{\mathrm{FM}}$, $\lambda_{\mathrm{bij}}$, $\lambda_{\mathrm{orth}}$) independently.
Appendix F Comprehensive results for vision-language model combinations
To provide a comprehensive evaluation of vision–language alignment, we test our proposed model across multiple vision and language model combinations on two benchmark datasets, CIFAR-10 and CINIC-10. The vision models considered include CLIP [Ramesh et al., 2022], ConvNeXt [Liu et al., 2022b], DINO [Caron et al., 2021], DINOv2 [Oquab et al., 2023], and DeiT [Touvron et al., 2021]. For the language models, we include the RN50x4 model from CLIP [Ramesh et al., 2022] as well as three models from the SentenceTransformers library [Reimers and Gurevych, 2019]: all-MiniLM-L6-v2, all-mpnet-base-v2, and all-Roberta-large-v1.
The results are summarized in Figure 5, where the error bars indicate the standard deviation computed over 20 random seeds.
Figure 5: Vision–language accuracy of the proposed model on combinations of multiple vision models with four language models on CIFAR-10 (top row) and CINIC-10 (bottom row)
These results indicate that our proposed approach consistently achieves higher matching accuracies across diverse vision–language encoder combinations, outperforming state-of-the-art baselines such as the Hahn-Grant solver in most configurations (cf. Figure 4 in the Hahn-Grant paper [Schnaus et al., 2025]).
Our method shows particularly strong performance with DINO and DINOv2 models, where most configurations achieve matching accuracies above 0.8 on both CIFAR-10 and CINIC-10 datasets. The CLIP models also demonstrate competitive performance, with several variants reaching near-perfect accuracy. Several models achieve perfect accuracy on CIFAR-10, including CLIP: RN50x4 and ViT-B/16, all ConvNeXt variants except CN-B-22, most DeiT models, and all DINO models. Comparable trends are observed on CINIC-10, where numerous models also reach 100% accuracy. The results indicate that the choice of pre-training model has a greater influence on performance than model size. DINO models exhibit remarkable consistency, achieving near-perfect accuracy in most configurations. In contrast, some larger models, such as CLIP: RN101 and RN50x16, perform poorly (33–40% on CIFAR-10), indicating that model scale alone does not guarantee superior performance.
Among the different language encoders, the sentence-transformer models (all-MiniLM-L6-v2, all-mpnet-base-v2, and All-Roberta-large-v1) outperform RN50x4, as they are specifically optimized for semantic text representation and generating high-quality text embeddings. In contrast, the RN50x4 encoder in CLIP is trained with an objective that prioritizes vision–language alignment rather than producing rich text embeddings.
We further evaluate the proposed model across diverse vision–language combinations on the larger-scale CIFAR-100 and ImageNet-100 benchmarks. Table 7 summarizes the results.These experimental results demonstrate that model performance is highly context-dependent, with no single architecture achieving universal superiority across CIFAR-100 and ImageNet-100.
Table 7: Vision-language alignment accuracies on CIFAR-100 and ImageNet-100 with two language models.
Model CIFAR-100 ImageNet-100
all-mpnet-base-v2 All-Roberta-large-v1 all-mpnet-base-v2 All-Roberta-large-v1
CLIP
RN50x16 46.67 ± 2.87 67.00 ± 4.97 40.67 ± 0.94 63.67 ± 3.40
RN50x64 48.33 ± 0.47 76.33 ± 0.94 37.33 ± 0.47 58.00 ± 4.55
ViT-L/14 46.00 ± 7.87 85.67 ± 1.89 74.00 ± 0.82 61.00 ± 1.41
ViT-L/14@336 79.67 ± 2.05 81.00 ± 1.41 41.33 ± 8.26 45.33 ± 14.20
DeiT
DeiT-B/16 47.00 ± 0.82 84.00 ± 0.82 67.33 ± 0.47 58.67 ± 4.03
DeiT-B/16@384 54.67 ± 5.44 88.00 ± 0.00 35.33 ± 0.47 57.67 ± 7.54
DeiT-B/16d 58.33 ± 8.63 58.67 ± 15.22 38.33 ± 6.02 54.33 ± 7.04
DeiT-B/16d@384 47.33 ± 8.99 42.67 ± 6.34 67.33 ± 4.50 65.67 ± 0.94
DINOv2
ViT-B/14 48.33 ± 0.47 55.00 ± 6.53 83.67 ± 0.47 60.33 ± 0.94
ViT-S/14 48.67 ± 4.64 58.33 ± 3.22 35.67 ± 0.47 63.67 ± 4.50
ViT-L/14 79.67 ± 1.70 60.67 ± 5.19 48.33 ± 2.49 69.67 ± 7.13
ViT-G/14 67.67 ± 1.70 58.33 ± 0.47 44.33 ± 8.26 49.33 ± 0.47